# CAD-Tokenizer: Towards Text-based CAD Prototyping via Modality-Specific Tokenization

### v1

@@ Abstract
Computer-Aided Design (CAD) is a foundational component of industrial prototyping. where models are defined not by raw coordinates but by construction sequences such as sketches and extrusions. This sequential structure enables both efficient prototype initialization and subsequent editing. Text-guided CAD prototyping, which unifies Text-to-CAD generation and CAD editing, has the potential to streamline the entire design pipeline. However, prior work has not explored this setting, largely because standard large language model (LLM) tokenizers decompose CAD sequences into natural-language word pieces, failing to capture primitive-level CAD semantics and hindering attention modules from modeling geometric structure. We conjecture that a multimodal tokenization strategy, aligned with CAD’s primitive and structural nature, can provide more effective representations. To this end, we propose CAD-Tokenizer, a framework that represents CAD data with modality-specific tokens using a sequence-based VQ-VAE with primitive-level pooling and constrained decoding. This design produces compact, primitive-aware representations that align with CAD’s structural nature. Applied to unified text-guided CAD prototyping, CAD-Tokenizer significantly improves instruction following and generation quality, achieving better quantitative and qualitative performance over both general-purpose LLMs and task-specific baselines.
@@ 1 Introduction
Recent work has begun to explore text-guided CAD subtasks, including CAD editing (Yuan et al., 2025) and text-to-CAD generation (Wang et al., 2025; Li et al., 2025b; Khan et al., 2024b). These studies not only demonstrated the feasibility of each individual task but also showed that large language models (LLMs) can effectively operate on CAD sequences. However, real-world design is rarely a one-shot process: engineers iteratively refine prototypes by alternating between creation and modification. An effective text-based CAD system should therefore support both tasks in a unified manner—generating CAD objects from instructions while also allowing editing of existing designs.
@@
    We leverage the formal grammar of CAD by introducing a finite-state automaton–based sampling strategy, which enforces valid syntax in generation and improves quality.
@@ 2.1 CAD Modeling
Classical CAD Tasks. Classical CAD tasks focus on generating CAD sequences from non-textual inputs. This includes hand sketches (Seff et al., 2022), point clouds (Khan et al., 2024a; Ma et al., 2024), sequence prefixes (Xu et al., 2022; 2023), or even random noise (Wu et al., 2021). There are also tasks involving masked variation, where some primitives are masked out and converted into random variants (Zhang et al., 2025; Xu et al., 2022; 2023). Our work differs from these works in two ways. First, our work focuses on CAD generation from textual inputs. Second, while works in these categories often use transformer encoders to extract features and decode specific IDs, they do not involve texts or natural languages and thus have no need to leverage the textual understanding capability of LLMs and hence do not need to align these IDs with them.
Text-based CAD Tasks. Text-based CAD tasks focus on generating CAD sequences based on textual instructions. Inputs for these tasks fall into two main types: the first type includes only textual instructions on how to generate an initial CAD model, and the second type includes a CAD sequence (of the same representation format) and a textual instruction on how to modify the given model. The former is the Text-to-CAD task (Wang et al., 2025; Li et al., 2024; 2025b; Khan et al., 2024b) and the latter is the CAD-editing task (Yuan et al., 2025). Our work differs from these tasks as we propose the text-based CAD prototyping task, which merges them. Given that initiating and editing CAD models is a standard procedure in industry, a single-model solution is both necessary and promising for the prototyping pipeline.
@@ 2.2 LLM Tokenization
LLMs operate over discrete tokens that are mapped into hidden embeddings. In natural language, these tokens are defined by vocabularies (Pennington et al., 2014; Mikolov et al., 2013) or obtained via algorithms such as byte-pair encoding (BPE) (Gage, 1994). In multimodal tasks where inputs often lie outside the space of predefined vocabularies, external encoders and decoders such as VAEs and VQ-VAEs (Kingma & Welling, 2022; van den Oord et al., 2018), vision transformers (Dosovitskiy et al., 2021; Radford et al., 2021), or diffusion-based algorithms (Podell et al., 2023) are commonly used to map continuous modalities into token sequences that LLMs can process.
@@
Prior CAD works are task-specific on tokenizations. Classical CAD generation methods train separate encoders and decoders for different inputs, such as sketches and extrusions (Xu et al., 2022; 2023; Khan et al., 2024b; a), resulting in multi-encoder pipelines. Recent text-based CAD efforts typically rely on off-the-shelf LLM tokenizers, directly applying LLM tokenizations to CAD sequences without adaptation (Wang et al., 2025; Yuan et al., 2025; Zhang et al., 2025). In contrast, our work introduces the first encoder–decoder tokenizer specifically designed to integrate with LLMs. We employ a single VQ-VAE–variant encoder–decoder pair to tokenize all CAD inputs and align its vocabulary with the LLM backbone. This unified tokenization allows the model to reason over CAD primitives as discrete units, addressing the fragmentation issues inherent in natural-language tokenizers.
@@ 3.2 Primitive-based VQ-VAE as CAD Tokenizer
where EMD is the squared Earth Mover’s Distance Loss, and $\overline{T}^{E}_{j}$ denotes the pooled vector of the tokens construct up the $j$-th primitive. We demonstrate this method in Figure 2(1) and provide additional details such as the detailed pooling procedure in Appendix B.
Aligning CAD Tokenizer with LLM Embeddings.
The raw primitive vectors $\{p^{\prime}_{j}\}$ have dimension $d_{vq}$, while the LLM requires tokens in its embedding space of dimension $d_{tok}$. Typically, this adaptation is performed during co-training with LLM backbones (Liu et al., 2023), which is compute-intensive and does not guarantee alignment in the reverse direction. We introduce a novel use of adapters that map VQ-VAE outputs and LLM embeddings bi-directionally. In detail, we leverage the frozen embedding and logit layers of the pretrained LLM.
@@
$\mathcal{L}_{\text{recon}}=\sum_{j\in k}\|\hat{p}^{\prime}_{j}-p^{\prime}_{j}\|_{2}^{2}.$
$\hat{p}^{\prime}_{j}=W^{d_{tok}}_{d_{vq}}(L_{embed}(p_{j}))$ is the reconstructed embedding corresponding to $\{p^{\prime}_{j}\}$. This alignment produces native LLM-recognizable primitive IDs, enabling seamless integration without modifying the backbone. The alignment procedure is illustrated by Figure 2(2).
@@ 3.4 Model Inference and Token Decoding
At each step, the FSA provides a series of masks restricting the candidate logits to grammar-compliant tokens, ensuring syntactic validity. The automaton state evolves with the decoder, whose choice for the next token is also passed to the automaton for deciding if a transition is going to take place. Despite some context-dependent geometric constraints are unable to fully catched, this method significantly reduces syntactic errors and improves sampling robustness by preventing obvious grammatical mistakes. Further implementation details are given in Appendix B.2 and the FSA design is illustrated in Figure 8.
@@ 4.1 Experimental Setups
Datasets. For training the VQ-VAE, we convert the SkexGen (Xu et al., 2022) dataset into the formatting used by CADFusion (Wang et al., 2025) and CAD-Editor (Yuan et al., 2025). We include the formatting details in the Appendix A. To prevent a distribution gap for the tokenizer between seen and unseen data, we use half of the training data for VQ-VAE training, ensuring the LLM also encounters samples not seen by the VQ-VAE encoder during its training. For training the backbone LLM, we use the datasets of CADFusion and CAD-Editor. Although both originate from SkexGen, the two datasets have different content, so we merge them for training our model without further deduplication. Since CAD-Editor uses the original split of SkexGen but the entire CADFusion dataset is split from its training set, we consolidate the CADFusion dataset and additionally annotated 200 pairs for our test set. We also balance this combined dataset by randomly selecting five CAD-Editing data points for every Text-to-CAD data point. Ultimately, this results in approximately 100k training pairs and 1000 testing pairs.
Implementation Details. The CAD sequence formalization and primitive selection is detailed in Appendix A. The VQ-VAE is trained with 5 encoder and 5 decoder layers, a codebook size of 2048, and an internal dimension of 512. It is trained with a learning rate of 3e-4 for up to 250 epochs, using early stopping with a patience of 25. The logit and embedding layers inserted are from LLaMA-3-8b-Instruct (4096-dim), which is also used as our LLM backbone. The alignment uses exactly the same setup, but the encoder, decoder and LLM layers are frozen, and only the two adapter layers are trained. For LLM training, the backbone is configured with a maximum token length of 1024 and a hidden dimension of 4096. We adopt Low-Rank Adaptation (LoRA) (Hu et al., 2021) for efficient fine-tuning, with a rank of 32 and an $\alpha$ value of 32. The model is trained for 20 epochs with a learning rate of 1e-4 and the AdamW optimizer. The VQ-VAE training is run on a single NVIDIA A100-80G GPU, and the LLM fine-tuning is conducted on four NVIDIA A6000-48GB SMX GPUs using Distributed Data Parallel (DDP).
@@
Metrics. Our evaluation focuses on different aspects of the generated CAD models. First, we measure the pairwise correspondence between generated outputs and the ground truth sequences. For this criterion, we use F1 scores and Chamfer Distance (CD). We follow CADFusion’s (Wang et al., 2025) measurements of F1-Sketch and F1-Extrusion for simplicity. Second, we measure distributional similarity. Between the generated samples and the ground truths, we follow (Xu et al., 2022; Yuan et al., 2025; Wang et al., 2025) and measure: Coverage (COV), which quantifies how well the generated distribution covers the ground truth; Minimum Matching Distance (MMD), which evaluates similarity by finding the closest matches between samples from the two distributions; and Jensen-Shannon Divergence (JSD) for distributional similarity. Third, we measure the Invalidity Ratio (IR) as an indicator of sequence robustness; and fourth, we measure the generated outputs’ alignment with instructions using a VLM-based score (VLM) and Human Evaluation (HE). The VLM is prompted to rate instruction following on a scale of 0 to 10, and human judges are asked to rank the outputs from different models based on the same criterion. Details on VLM prompting and human evaluation methodology can be found in Appendix C.1.
@@ 4.2 Main Results
Table 2: The reconstruction quality of the CAD-Tokenizer variants and HNC-CAD. Only the sketch scores are reported for the F1 score because the CAD-Tokenizer encodes objects by sketch-extrusion pairs and always achieves full score for this sub-metric. CAD-Tokenizer (curve) is the default variant which we reported in the main quantitative results.
@@ 4.3 Ablation Studies
Table 3: The performance of each tokenizer when integrated and trained with the backbone LLM. CAD-Tokenizer (curve) is the default variant which we reported in the main quantitative results, while loop is a variant on pooling loop primitives and single does not perform specific pooling.
@@
Table 4: The generation quality of each sampling method after training. The base model used for evaluation is the fine-tuned CAD-Tokenizer. CAD-Tokenizer(+FSA) is the default variant which we reported in the main quantitative results.
@@ 5 Limitations
Our work’s limitations are two-fold. First, the quality gap between the open-source and private-sector CAD data limits us from training on further complex shapes. Second, as mentioned and referred, we observe a gap between the distributional metrics and the actual performance in CAD Editing, which stems from the failure to evaluate the ’intent to keep the original shape.’ We therefore anticipate both better datasets and metrics that can create a better ground for future works to improve. Additionally, our failure study in Appendix C.2.3 suggests better spatial or commonsense reasoning capabilities, which can only be improved by better pretrained backbones. We include some of the related discussions, as well as the detail of our failure study, in Appendix C.2.
@@ A.1 Sequence Representation
We follow the practice of Zhang et al. (2025); Yuan et al. (2025); Wang et al. (2025). In detail, a CAD object is defined by a series of sketches and extrusions, termed the Sketch-and-Extrude Modeling (SEM) format.
@@ B.2 FSA Sampling
Figure 8: An overview of our FSA design. At each step, the FSA receives an input action, transitions to the corresponding new state (i.e., updates its internal state), and returns the mask(s) associated with that new state (node).
Algorithm 1 The FSA-driven decoding algorithm.
 Given: LLM-generated primitive tokens $T$
 Given: Max length of generation max_lenmax\_len
$State\leftarrow\texttt{`init'}$
$M\leftarrow[\texttt{initial\_mask}]$
$Seq\leftarrow\emptyset$
while |Seq|<max_len|Seq|<max\_len do
  $m\leftarrow$ $M$.next()
  logits $\leftarrow$ Decoder(SeqSeq, $T$)
[…]
  SeqSeq.insert(choice)
  if $M=\emptyset$ then
   new_state, new_M = FSA(StateState, choice)
   $State\leftarrow$ new_state
   $M\leftarrow$ new_M
  end if
end while

### v2

@@ Abstract
Computer-Aided Design (CAD) is a foundational component of industrial prototyping, where models are defined not by raw coordinates but by construction sequences such as sketches and extrusions. This sequential structure enables both efficient prototype initialization and subsequent editing. Text-guided CAD prototyping, which unifies Text-to-CAD generation and CAD editing, has the potential to streamline the entire design pipeline. However, prior work has not explored this setting, largely because standard large language model (LLM) tokenizers decompose CAD sequences into natural-language word pieces, failing to capture primitive-level CAD semantics and hindering attention modules from modeling geometric structure. We conjecture that a multimodal tokenization strategy, aligned with CAD’s primitive and structural nature, can provide more effective representations. To this end, we propose CAD-Tokenizer, a framework that represents CAD data with modality-specific tokens using a sequence-based VQ-VAE with primitive-level pooling and constrained decoding. This design produces compact, primitive-aware representations that align with CAD’s structural nature. Applied to unified text-guided CAD prototyping, CAD-Tokenizer significantly improves instruction following and generation quality, achieving better quantitative and qualitative performance over both general-purpose LLMs and task-specific baselines.
@@ 1 Introduction
Recent work has begun to explore text-guided CAD subtasks, including CAD editing (Yuan et al., 2025) and text-to-CAD generation (Wang et al., 2025a; Li et al., 2025b; Khan et al., 2024b). These studies not only demonstrated the feasibility of each individual task but also showed that large language models (LLMs) can effectively operate on CAD sequences. However, real-world design is rarely a one-shot process: engineers iteratively refine prototypes by alternating between creation and modification. An effective text-based CAD system should therefore support both tasks in a unified manner—generating CAD objects from instructions while also allowing editing of existing designs.
@@
    We leverage the formal grammar of CAD by introducing a finite-state automaton (FSA)–based sampling strategy, which enforces valid syntax in generation and improves quality.
@@ 2.1 CAD Modeling
Classical CAD Tasks. Classical CAD tasks focus on generating CAD sequences from non-textual inputs. This includes hand sketches (Seff et al., 2022; Xu et al., 2025), point clouds (Khan et al., 2024a; Ma et al., 2024), sequence prefixes (Xu et al., 2022; 2023), or even random noise (Wu et al., 2021). There are also tasks involving masked variation, where some primitives are masked out and converted into random variants (Zhang et al., 2025; Xu et al., 2022; 2023). Our work differs from these works in different ways. First, our work focuses on CAD generation from textual inputs. Second, while works in these categories often use transformer encoders to extract features and decode specific IDs, they do not unify the encoding procedure into one model due to technical limitations. Third, while some of the previous works (Xu et al., 2022; 2023) train transformer modules using their codebooks, they do not involve texts or natural language inputs and thus have no need to leverage pretrained LLMs. This prevents the need to add alignment modules in the classic paradigms.
Text-based CAD Tasks. Text-based CAD tasks focus on generating CAD sequences based on textual instructions. Inputs for these tasks fall into two main types: the first type includes only textual instructions on how to generate an initial CAD model, and the second type includes a CAD sequence (of the same representation format) and a textual instruction on how to modify the given model. The former is the Text-to-CAD task (Wang et al., 2025a; Li et al., 2024; 2025b; Khan et al., 2024b; Govindarajan et al., 2025; Guan et al., 2025; Wang et al., 2025b) and the latter is the CAD-editing task (Yuan et al., 2025). Our work differs from these tasks as we propose the text-based CAD prototyping task, which merges them. Given that initiating and editing CAD models is a standard procedure in industry, a single-model solution is both necessary and promising for the prototyping pipeline.
@@ 2.2 LLM Tokenization
LLMs operate over discrete tokens that are mapped into hidden embeddings. In natural language, these tokens are defined by vocabularies (Pennington et al., 2014; Mikolov et al., 2013) or obtained via algorithms such as byte-pair encoding (BPE) (Gage, 1994). In multimodal tasks where inputs often lie outside the space of predefined vocabularies, external encoders and decoders such as VAEs and VQ-VAEs (Kingma and Welling, 2022; van den Oord et al., 2018), vision transformers (Dosovitskiy et al., 2021; Radford et al., 2021), or diffusion-based algorithms (Podell et al., 2023) are commonly used to map continuous modalities into token sequences that LLMs can process.
@@
Prior CAD works are task-specific on tokenizations. Classical CAD generation methods train separate encoders and decoders for different inputs, such as sketches and extrusions (Xu et al., 2022; 2023; Khan et al., 2024b; a), resulting in multi-encoder pipelines. Recent text-based CAD efforts typically rely on off-the-shelf LLM tokenizers, directly applying LLM tokenizations to CAD sequences without adaptation (Wang et al., 2025a; Yuan et al., 2025; Zhang et al., 2025). In contrast, our work introduces the first encoder–decoder tokenizer specifically designed to integrate with LLMs. We employ a single VQ-VAE–variant encoder–decoder pair to tokenize all CAD inputs and align its vocabulary with the LLM backbone. This unified tokenization allows the model to reason over CAD primitives as discrete units, addressing the fragmentation issues inherent in natural-language tokenizers.
@@ 3.2 Primitive-based VQ-VAE as CAD Tokenizer
where EMD is the squared Earth Mover’s Distance Loss, $\{p^{\prime}\}$ denotes the set of all primitive latent vectors $\{p^{\prime}_{1},...,p^{\prime}_{k}\}$ generated by the encoder and pooling layer, $\text{1}_{i}$ denotes the input data’s one-hot property, and $\overline{T}^{E}_{j}$ denotes the pooled vector of the tokens construct up the $j$-th primitive. We demonstrate this method in Figure 2(1) and provide additional details such as the detailed pooling procedure in Appendix B.
Aligning CAD Tokenizer with LLM Embeddings. The raw primitive vectors $\{p^{\prime}_{j}\}$ have dimension $d_{vq}$, while the LLM requires tokens in its embedding space of dimension $d_{tok}$. Typically, this adaptation is performed during co-training with LLM backbones (Liu et al., 2023), which is compute-intensive and does not guarantee alignment in the reverse direction. We introduce a novel use of adapters that map VQ-VAE outputs and LLM embeddings bi-directionally. In detail, we leverage the frozen embedding and logit layers of the pretrained LLM.
@@
[…]
$\hat{p}^{\prime}_{j}=W^{d_{tok}}_{d_{vq}}(L_{embed}(p_{j}))$ is the reconstructed embedding corresponding to $\{p^{\prime}_{j}\}$. The latter term is the reconstruction of the post-embedding layer tokens and the pre-logit tokens to ensure both adapter layers are trained. This alignment produces native LLM-recognizable primitive IDs, enabling seamless integration without modifying the backbone. The procedure is illustrated by Figure 2(2).
@@ 3.4 Model Inference and Token Decoding
At each step, the FSA provides a series of masks restricting the candidate logits to grammar-compliant tokens, ensuring syntactic validity. The automaton state evolves with the decoder, whose choice for the next token is also passed to the automaton for deciding if a transition is going to take place. Despite some context-dependent geometric constraints are unable to fully catched, this method significantly reduces syntactic errors and improves sampling robustness by preventing obvious grammatical mistakes. We demonstrate the detailed process by Algorithm 1 and the FSA design in Figure 3. Further implementation details are given in Appendix B.2.
Figure 3: An overview of our FSA design. At each step, the FSA receives an input action, transitions to the corresponding new state (i.e., updates its internal state), and returns the mask(s) associated with that new state (node). The FSA map is specific to our CAD sequence representation, which can be found in Appendix A.
Algorithm 1 The FSA-driven decoding algorithm.
 Given: LLM-generated primitive tokens $T$
 Given: Max length of generation max_lenmax\_len
$State\leftarrow\texttt{`init'}$
$M\leftarrow[\texttt{initial\_mask}]$
$Seq\leftarrow\emptyset$
while |Seq|<max_len|Seq|<max\_len do
  $m\leftarrow$ $M$.next()
  logits $\leftarrow$ Decoder(SeqSeq, $T$)
[…]
  SeqSeq.append(choice)
  if $M=\emptyset$ then
   new_state, new_M = FSA(StateState, choice)
   $State\leftarrow$ new_state
   $M\leftarrow$ new_M
  end if
end while
@@ 4.1 Experimental Setups
Datasets. For training the VQ-VAE, we convert the SkexGen (Xu et al., 2022) dataset into the formatting used by CADFusion (Wang et al., 2025a) and CAD-Editor (Yuan et al., 2025). We include the formatting details in the Appendix A. To prevent a distribution gap for the tokenizer between seen and unseen data, we use half of the training data for VQ-VAE training, ensuring the LLM also encounters samples not seen by the VQ-VAE encoder during its training. For training the backbone LLM, we use the datasets of CADFusion and CAD-Editor. Although both originate from SkexGen, the two datasets have different content, so we merge them for training our model without further deduplication. Since CAD-Editor uses the original split of SkexGen but the entire CADFusion dataset is split from its training set, we consolidate the CADFusion dataset and additionally annotated 200 pairs for our test set. We also balance this combined dataset by randomly selecting five CAD-Editing data points for every Text-to-CAD data point. Ultimately, this results in approximately 100k training pairs and 1000 testing pairs.
Implementation Details. The CAD sequence formalization and primitive selection is detailed in Appendix A. The VQ-VAE is trained with 5 encoder and 5 decoder layers, a codebook size of 2048, and an internal dimension of 512. It is trained with a learning rate of 3e-4 for up to 250 epochs, using early stopping with a patience of 25. The logit and embedding layers inserted are from LLaMA-3-8b (4096-dim), which is also used as our LLM backbone. The alignment uses exactly the same setup, but the encoder, decoder and LLM layers are frozen, and only the two adapter layers are trained. For LLM training, the backbone is configured with a maximum token length of 1024 and a hidden dimension of 4096. We adopt Low-Rank Adaptation (LoRA) (Hu et al., 2021) for efficient fine-tuning, with a rank of 32 and an $\alpha$ value of 32. The model is trained for 20 epochs with a learning rate of 1e-4 and the AdamW optimizer. The VQ-VAE training is run on a single NVIDIA A100-80G GPU, and the LLM fine-tuning is run by four NVIDIA A6000-48GB SMX GPUs using Distributed Data Parallel (DDP).
@@
Metrics. Our evaluation focuses on different aspects of the generated CAD models. First, we measure the pairwise correspondence between generated outputs and the ground truth sequences. For this criterion, we use F1 scores and Chamfer Distance (CD). We follow CADFusion’s (Wang et al., 2025a) measurements of F1-Sketch and F1-Extrusion for simplicity. Second, we measure distributional similarity. Between the generated samples and the ground truths, we follow (Xu et al., 2022; Yuan et al., 2025; Wang et al., 2025a) and measure: Coverage (COV), which quantifies how well the generated distribution covers the ground truth; Minimum Matching Distance (MMD), which evaluates similarity by finding the closest matches between samples from the two distributions; and Jensen-Shannon Divergence (JSD) for distributional similarity. Third, we measure the Invalidity Ratio (IR) as an indicator of sequence robustness; and fourth, we measure the generated outputs’ alignment with instructions using a VLM-based score (VLM) and Human Evaluation (HE). The VLM is prompted to rate instruction following on a scale of 0 to 10, and human judges are asked to rank the outputs from different models based on the same criterion. Details on VLM prompting and human evaluation methodology can be found in Appendix C.1.
@@ 4.2 Main Results
Table 2: The ablation study on the reconstruction quality of the CAD-Tokenizer variants and HNC-CAD. Only the sketch scores are reported for the F1 score because the CAD-Tokenizer encodes objects by sketch-extrusion pairs and always achieves full score for this sub-metric. CAD-Tokenizer (curve) is the default variant that we reported in main results.
@@ 4.3 Ablation Studies
Table 3: The ablation study of each tokenizer when integrated and trained with the backbone LLM. CAD-Tokenizer (curve) is the default variant which we reported in the main quantitative results, while loop is a variant on pooling loop primitives and single does not perform specific pooling.
@@
Table 4: The ablation study on the generation quality of each sampling method after training. The base model used for evaluation is the fine-tuned CAD-Tokenizer. CAD-Tokenizer(+FSA) is the default variant which we reported in the main quantitative results.
@@ 5 Limitations
Our work’s limitations are two-fold. First, the quality gap between the open-source and private-sector CAD data limits us from training on further complex shapes. Second, as mentioned and referred, we observe a gap between the distributional metrics and the actual performance in CAD Editing, which stems from the failure to evaluate the ‘intent to keep the original shape.’ We therefore anticipate both better datasets and metrics that can create a better ground for future works to improve. Additionally, our failure study in Appendix C.2.3 suggests better spatial or commonsense reasoning capabilities, which can only be improved by better pretrained backbones. We include some of the related discussions, as well as the detail of our failure study, in Appendix C.2.
@@ 6 Conclusion
The Use of Large Language Models
The use of LLMs is restricted in revising the word choices and grammar only. No LLMs are used for research ideation or contributive enough to be regarded as a contributor.
Acknowledgements
We would like to thank Lexin Zhou, Rushuai Yang, Yun Pan for insightful discussions. We appreciate the discussion with Zhuoyu Jiang and Yuheng Lu during the rebuttal phase. We also acknowledge the support from the MSR Asia Industry Innovation Center for motivating real-world examples and guidance on industrial pipelines.
@@ A.1 Sequence Representation
We follow the practice of Zhang et al. (2025); Yuan et al. (2025); Wang et al. (2025a). In detail, a CAD object is defined by a series of sketches and extrusions, termed the Sketch-and-Extrude Modeling (SEM) format.
@@ B.2 FSA Sampling
  Methods COV JSD MMD VLM(left)
GPT-4o <–> Original Seq. 60.8 6.75 2.39 1.94
CAD-Tokenizer <–> Original Seq. 52.4 2.54 1.97 5.09
Original Seq. <–> Ground Truth 55.6 6.07 2.54 2.02